# Seeing Through Green: Text-Based Classification and the Firm’s Returns from Green Patents

### v1

@@ A.1. Green-patents vocabulary
Table A1: Baseline key-expresions.
Baseline keywords
alternative energy waste management energy recovery ozone generators
anaerobic digestion waste reduction energy savings photocatalysis
carbon capture water conservation engine optimization plant based
carbon footprint aerodynamic design environmental certification pollution control
carbon sequestration air purification environmental compliance precision agriculture
circular economy bio-based materials environmental conservation product lifespan
clean energy biodegradable polymers environmental labeling public transportation
clean production biodegradable products environmental management rainwater harvesting
composting biodiversity environmental policy recyclable materials
conservation biofuel environmental stewardship recycled materials
dioxide capture biomass fuel cells recycled packaging
energy conservation bioplastics geothermal recycled products
energy efficient biotechnology green buildings resource efficient
energy optimization carbon filters green design smart grids
environmental impact carbon neutral green innovation solar geoengineering
environmental risk carbon offsetting green manufacturing solar panels
geothermal power cellulosic plastics green technology sustainability metrics
global warming clean technology green transportation sustainability practices
green chemistry climate adaptation greenhouse gas sustainable agriculture
hydropower climate friendly greenhouse gases sustainable farming
low carbon compostable materials hybrid vehicles sustainable forestry
low emission eco conscious hydroelectric dams sustainable manufacturing
membrane filtration eco efficient hydroponics sustainable materials
membrane separation eco friendly insulation sustainable practices
recycling eco innovation leed certified vertical farming
regenerative braking eco solutions life cycle waste minimization
renewable energy ecodesign lightweight materials waste recycling
resource recovery ecological design low impact water efficient
reverse osmosis ecological footprint material compatibility water treatment
sustainable electric motors natural ventilation wind turbines
sustainable development electric vehicles organic farming zero emissions
urban planning energy efficiency ozone depletion zero waste
Table A2: Keywords for single REGEX match
Expressions for single match
aerogenerator carbon filters ecological design geothermal photocatalysis soilless
aerodynamic design carbon neutral ecological footprint geothermal power photocatalyst soilless cultivation
aeroponic carbon offset ecological restoration geothermy photocatalytic soilless culture
afforestation carbon offsetting ecologically global warming photodegradable solar geoengineering
air purification carbon sequestration ecology green buildings photolysis solar panels
alternative energy carbonfilter ecosystem green chemistry photolytic solarenergy
anaerobic catalyzing electric motors green design photooxidation solarpanel
anaerobic digestion cellulosic plastics electric powertrain green innovation photovoltaic solarpower
bioactivity circular economy electric vehicle green manufacturing plant based sustainability
bio-based materials clean energy electric vehicles green technology pollution control sustainability metrics
biocarbon clean production electricvehicle green transportation polylactide sustainability practices
biochemistry clean technology emission reduction greenhouse gas precision agriculture sustainable
biocomposite climate adaptation energy conservation greenhouse gases product lifespan sustainable agriculture
biodegradable climate friendly energy efficiency greening public transportation sustainable development
biodegradable compostable compostable energy efficient hybrid powertrain rainwater harvesting sustainable farming
biodegradable polymer compostable material energy optimization hybrid vehicle recyclable sustainable forestry
biodegradable polymers compostable materials energy recovery hybrid vehicles recyclable material sustainable manufacturing
biodegradable product composted energy save hydroelectric recyclable materials sustainable materials
biodegradable products composting energy saving hydroelectric dam recyclate sustainable practices
biodegrade conservancy hydropower energy savings hydroelectric dams recycle material vertical farming
biodiesel conservation energysaving hydroponics recycle practice thrift waste minimization
biodiversity decarbonation engine optimization hydropower recycle product waste recycle
bioenergy decarbonization environmental certification insulation recycleable waste recycling
bioethanol dechlorinate environmental compliance leed certified recycled waste reduction
biofertilizer decyanation environmental conservation life cycle recycled materials water conservation
biofuel demercuration environmental labeling lightweight materials recycled packaging water efficient
bioleache depollute environmental management lignocellulose recycled products water purification
biologically degradable depollution environmental policy low emission recycling water treatment
biomass development environmental stewardship low impact regenerative braking waterpower
bioplastic dioxide capture environmentally friendly mangrove forest renewable energy wind turbines
bioplastics eco conscious environmentally sustainable material compatibility resource efficient windmill
biopolymer eco efficient fertigation natural ventilation resource recovery windpower
bioproduct eco friendly footprint accounting organic farming rooftop greening windturbine
bioremediation eco innovation forward osmosis oxyfuel seawater desalination windwheel
biotechnology eco solutions fuel cells ozone depletion seawater desalting zero emissions
carbon capture ecodesign fuelcell ozone generators smart grids zero waste
  • Note: The table reports the list of keywords for single REGEX match into patents texts, i.e. a text containing at least one of the expression, if classified as “green” by code-based algorithms, were classified as “true green”.
Table A3: Keywords combinations for co-occurrence REGEX match (part I).
Keyword Co-occurrence
activatedcarbon Filtration OR purification OR pollutants
aeroponic cultivation Hydroponics OR vertical farming
aeroponic culture Hydroponics OR vertical farming
bamboo charcoal Sustainable forestry
biochemical Biofuels OR bioplastics
bioconversion Biomass OR renewable energy
butanol Biofuels
carbon Capture
carwash Water-efficient
claus Sulfur recovery OR tail gas treatment
daylighting energy saving
deduste Air pollution
denitride Air pollution
denitrifying Water pollution
denitrogenation Air pollution
denox Air pollution
desalinate Energy-efficient
desalination Energy-efficient
desalinization Energy-efficient
desalt Energy-efficient
desulfurate Air pollution
disinfection Non-chlorine OR ultraviolet
disinfector Non-chlorine OR ultraviolet
dmfc Direct methanol fuel cell
electrochemical Fuel cell
electrodialysis Desalination OR wastewater treatment
electrolyzer Hydrogen production OR water electrolysis
environmental Sustainability OR pollution reduction
enzymolysis Biodegradable materials OR biomass conversion
ethanol Biofuel OR renewable energy
evaporative crystallization Water treatment OR desalination
fenton Wastewater treatment OR pollutant degradation
ferment Biofuels OR biogas production
fermentable sugar Biofuels OR biogas production
filter Air pollution control OR water treatment
flywheel Energy storage OR regenerative braking
greywater Water reuse OR wastewater treatment
humification Composting ganic matter
hydrolysed Biomass conversion OR biofuel production
hydrolytic acidification Biomass conversion OR biofuel production
hyperfiltration Water treatment OR desalination
insulate Energy efficiency OR building insulation
insulated Energy efficiency OR building insulation
insulating Energy efficiency OR building insulation
insulative Energy efficiency OR building insulation
insulator Energy efficiency OR building insulation
lignocellulosic Biomass conversion OR biofuel production
livestock excreta Biogas production OR waste management
livestock manure Biogas production OR waste management
membrane Filtration OR water treatment
methanation Biogas production OR renewable energy
methanogenesis Biogas production OR renewable energy
microfiltration Water treatment OR wastewater treatment
microgrid Renewable energy OR distributed generation
nanobubble Water treatment OR air purification
nanofilter Water treatment OR wastewater treatment
nanofiltration Water treatment OR wastewater treatment
nitrifying denitrify Wastewater treatment OR nitrogen removal
oilfield reinjection Enhanced oil recovery OR carbon capture
orc Organic Rankine Cycle OR waste heat recovery
cropping practice Sustainable OR organic
energysource Clean OR sustainable
foodwaste Reduction OR avoid
renewable Energy OR resource
  • Note: The table reports the list of keywords co-occurrences for REGEX match into patents texts, i.e. a text containing at least one of the expression combinations, if classified as “green” by code-based algorithms, were classified as “true green”.
Table A4: Keywords combinations for co-occurrence REGEX match (part II)
Keyword Co-occurrence
renewably Energy OR resource
reusable Material OR product OR resource
reuse Material OR product OR resource
ordered mesoporous catalysis
oxygenator Aquaculture OR wastewater treatment
oxygenerator Aquaculture OR wastewater treatment
ozonator Water treatment OR air purification
ozone Water treatment OR air purification
ozonize Water treatment OR air purification
ozonizer Water treatment OR air purification
ozonolysis Water treatment OR air purification
peak shaving energy management OR energy storage
pemfc Proton exchange membrane fuel cell OR hydrogen fuel cell
pervaporation Membrane separation OR desalination
photochemistry Solar energy conversion
photoreaction Solar energy conversion
photosynthetic Biofuels
poultry farming Sustainable agriculture OR waste management
poultry manure Sustainable agriculture OR waste management
propeller Electric OR hybrid OR low-emission
propulsion Electric OR hybrid OR low-emission
propulsor Electric OR hybrid OR low-emission
pyrohydrolysis Waste conversion OR biomass conversion
recuperation energy OR waste OR material
regasification Liquefied natural gas OR renewable energy sources
regenerative Braking OR energy recovery
remediation Environmental cleanup OR contaminated soil
reprocess Waste management OR resource recovery
reprocessed Waste management OR resource recovery
reprocessing Waste management OR resource recovery
resource Efficiency OR conservation
restoration Ecological restoration OR land reclamation
reusing Waste OR Material
reutilization Waste OR Material
reutilize Waste OR Material
reutilizing Waste OR Material
reverseosmosis Water desalination OR water treatment
rice husk Biomass conversion OR renewable energy source
saccharify enzyme Biofuel production OR biomass conversion
seedling raising Sustainable agriculture OR forestry
semicoke Biomass conversion OR coal alternative
simultaneously practice thrift Resource efficiency OR waste reduction
smart grid Renewable energy integration OR energy efficiency
sncr denitration Flue gas treatment OR air pollution control
sofc Solid oxide fuel cell OR clean energy
solar Energy OR cell OR panel
straw Biomass resource OR renewable energy
sugar beet Biofuel feedstock OR renewable energy
sun shading Sustainable design OR building efficiency
sunshade Sustainable design OR building efficiency
supercapacitor Energy storage OR renewable energy integration
supercharge Electric vehicle OR battery
supercharging Electric vehicle OR battery
thermocatalytic Biomass conversion OR waste valorization
thermochemical Biomass conversion OR renewable energy
ultrafilter Water treatment OR purification
ultrafiltrate Water treatment OR purification
ultrafiltration Water treatment OR purification
ultralight Material OR construction
ultralow Emission OR power consumption
uv photolysis Water treatment OR pollutant degradation
wasted Recycling OR resource recovery
wastepaper Recycling OR paper recycling
  • Note: The table reports the list of keywords co-occurrences for REGEX match into patents texts, i.e. a text containing at least one of the expression combinations, if classified as “green” by code-based algorithms, were classified as “true green”.
@@ A.3. Word2Vec specification
In this section, we provide details about the selection of the Word2Vec model’s hyperparameters. We need to find the context window size ($C$), which defines the context length for training the neural network, and the minimum count ($MC$), which sets the frequency threshold for a word to be included in the training corpus. We employ a data-driven tuning process. We use WordSim353, which is a test collection for measuring word similarity or relatedness (Agirre et al., 2009). We then train multiple Word2Vec models using a grid of different $W$ and $MC$ parameter combinations. For each trained model, we calculated the cosine similarity for word pairs in the WordSim353 list and measured the Pearson correlation between our results and the gold standard scores. As shown in Figure A1, all tested parameter pairs yielded a positive correlation above 50%. Generally, models with a smaller $C$ performed better when the minimum count was held constant. This empirical process pointed towards optimal values of approximately $MC$=40 and $C$=2, which we adopt in our analysis in the manuscript.
@@ B.1. Firm-level dataset distribution
01-03 Crop, animal production, fishing, hunting, forestry and logging 79,028 3.16
05-09 Mining and extraction 5,914 0.24
10-32 Manifacture 313,035 12.5
33 Repair and installation of machinery and equipment 22,318 0.89
35 Electricity, gas, steam and air conditioning supply 24,245 0.97
36-39 Water, sewerage and waste management 15,266 0.61
41-43 Construction, civil engineering and special activities 125,066 4.99
45-47 Wholesale and retail trade 630,142 25.16
49-51 Transport 97,219 3.88
52 Warehousing and support activities for transportation 26,388 1.05
53 Postal and courier activities 2,514 0.10
55 Accommodation 40,433 1.61
56 Food and beverage service activities 92,873 3.71
58 Publishing activities 14,375 0.57
59-63 ICT, motion pictures, television and broadcasting 96,546 3.86
64-68 Financial, insurance and real estate activities 284,155 11.35
69 Legal and accounting activities 76,168 3.04
70 Activities of head offices; management and consultancy 81,255 3.24
71 Architectural and engineering activities 76,801 3.07
72 Scientific research and development 6,824 0.27
73 Advertising and market research 28,086 1.12
74 Other professional, scientific and technical activities 37,452 1.50
75 Veterinary activities 5,314 0.21
77 Rental and leasing activities 18,562 0.74
78 Employment activities 6,942 0.28
79 Travel agency and tour operator 14,402 0.58
80 Security and investigation activities 6,712 0.27
81 Services to buildings and landscape activities 27,566 1.10
82 Office administrative and office support 34,011 1.36
@@
85 Education 32,992 1.32
86 Human health activities 74,445 2.97
87 Residential care activities 7,731 0.31
88 Social work activities without accommodation 7,793 0.31
90-91 Cultural and creative activities 13,935 0.56
92-93 Sports, amusement and gambling activities 25,394 1.01
94 Activities of membership organisations 1,340 0.05
95 Repair of computers and personal goods 10,743 0.43
96 Other personal service activities 39,633 1.58
97 Households as employers ofdomestic personnel 23 0.01
98 Undifferentiated goods andservices 7 0.01
99 Activities of extraterritorial organisations and bodies 51 0.01
@@
Austria 385 147 28
Belgium 529 179 44
Bulgaria 206 18 2
@@
Spain 1,200 232 41
Finland 813 213 44
France 2,681 668 170
Greece 16 3 1
Croatia 20 3 1
Hungary 108 25 4
@@
Netherlands 187 63 17
Poland 984 228 66
Portugal 125 30 4
Romania 146 44 20
Sweden 1,156 271 33
Slovenia 86 19 1
@@
01-03 Crop, animal production, fishing, hunting, forestry and logging 56 12 3

### v2

@@ A.1. Green-patents vocabulary
Table A2: Additional keywords for REGEX match (part I).
Keyword Co-occurrence Keyword Co-occurrence
acidogenesis wastewater, pollutant disinfector non-chlorine, ultraviolet
activatedcarbon filtration, purification, pollutants dmfc direct methanol fuel cell
adsorption emissions & capture ecofriendly -
aerogenerator - ecologic -
aeroponic hydroponics, vertical farming ecological restoration -
aeroponic cultivation hydroponics, vertical farming ecologically -
aeroponic culture hydroponics, vertical farming ecologization -
aeroponics cultivation, cropping ecology -
afforestation - ecosystem -
bamboo charcoal sustainable forestry ecotoxicity avoid, reduce, reduction
bioactivity - electric motor -
bioaugmentation - electric powertrain -
biocarbon - electric vehicle -
biochemical biofuels, bioplastics electricvehicle -
biochemistry biofuels, bioplastics electrochemical fuel cell
biocomposite - electrodialysis desalination, wastewater treatment
bioconversion biomass, renewable energy electrolyzer hydrogen production, water electrolysis
biodegradable - emission reduction -
biodegradable compostable - energy save -
biodegradable polymer - energy saving -
biodegradable product - energysaving -
biodegrade - energysource clean, sustainable, renewable
biodiesel - environmental sustainability, pollution reduction
bioenergy - environmental impact reduce, lower
bioethanol - environmental risk reduce, lower
biofertilizer - environmentally benign -
bioleache - environmentally friendly -
biologically degradable - environmentally sustainable -
bioplastic - enzymolysis biodegradable materials, biomass conversion
biopolymer - ethanol biofuel, renewable energy
bioproduct - evaporative crystallization water treatment, desalination
bioremediation - fenton wastewater treatment, pollutant degradation
butanol biofuels ferment biofuels, biogas production
carbon footprint reducing fermentable sugar biofuels, biogas production
carbon offset - fertigation -
carbonfilter - filter air pollution control, water treatment
carwash water-efficient flywheel energy storage, regenerative braking
catalyzing - foodwaste reduction, avoid
claus sulfur recovery, tail gas treatment footprint accounting -
co2 capture, reduction, reduce, remove, removal forward osmosis -
coemission avoid, reduction, reduce, remove, removal fuelcell -
compostable - geothermy -
compostable material - greening -
composted - greywater water reuse, wastewater treatment
conservancy hydropower - humification composting, organic matter
conservation minded - hybrid powertrain -
cropping practice sustainable, organic hybrid vehicle -
daylighting energy saving, energy efficient hydroelectric -
decarbonation - hydroelectric dam -
decarbonization - hydrolysed biomass conversion, biofuel production
dechlorinate - hydrolytic acidification biomass conversion, biofuel production
decyanation - hyperfiltration water treatment, desalination
deduste air pollution insulate energy efficiency, building insulation
demercuration - insulated energy efficiency, building insulation
denitride air pollution insulating energy efficiency, building insulation
denitrifying water pollution insulative energy efficiency, building insulation
denitrogenation air pollution insulator energy efficiency, building insulation
denox air pollution lignocellulose biomass conversion, biofuel production
depollute - lignocellulosic biomass conversion, biofuel production
depollution - livestock excreta biogas production, waste management
desalinate energy-efficient livestock manure biogas production, waste management
desalination energy-efficient low carbon emission
desalinization energy-efficient mangrove forest -
desalt energy-efficient membrane filtration, water treatment
desulfurate air pollution membrane filtration water treatment
disinfection non-chlorine, ultraviolet
  • Note: The table reports the list of keywords for REGEX match into patents texts produced with the W2V model, i.e. a text containing at least one of the expression combinations, if classified as “green” by standard classification algorithms, were classified as “true green”.
Table A3: Additional keywords for REGEX match (part II).
Keyword Co-occurrence Keyword Co-occurrence
membrane separation water treatment renewably energy, resource, material
methanation biogas production, renewable energy reprocess waste management, resource recovery, resource, waste, material
methanogenesis biogas production, renewable energy reprocessed waste management, resource recovery, resource, waste, material
microfiltration water treatment, wastewater treatment reprocessing waste management, resource recovery, resource, waste, material
microgrid renewable energy, distributed generation resource efficiency, conservation, reuse
nanobubble water treatment, air purification restoration ecological restoration, land reclamation
nanofilter water treatment, wastewater treatment reusable material, product, resource
nanofiltration water treatment, wastewater treatment reuse material, product, resource
natural ventilation energy efficient, energy saving reusing waste, material, resource
nitrifying denitrify wastewater treatment, nitrogen removal reutilization waste, material, resource
oilfield reinjection enhanced oil recovery, carbon capture reutilize waste, material, resource
orc organic rankine cycle, waste heat recovery reutilizing waste, material, resource
ordered mesoporous catalysis reverse osmosis low energy, sustainable, reuse, recycle, low power
oxyfuel - reverseosmosis low energy, sustainable, reuse, recycle, low power
oxygenator aquaculture, wastewater treatment rice husk biomass conversion, renewable energy source
oxygenerator aquaculture, wastewater treatment rooftop greening -
ozonator water treatment, air purification saccharify enzyme biofuel production, biomass conversion
ozone water treatment, air purification seawater desalination low energy, sustainable, reuse, recycle, low power
ozone depletion reduce, prevent seawater desalting low energy, sustainable, reuse, recycle, low power
ozone generators water treatment seedling raising sustainable agriculture, forestry
ozonize water treatment, air purification semicoke biomass conversion, coal alternative
ozonizer water treatment, air purification sequester carbon dioxide, co2, greenhouse gases
ozonolysis water treatment, air purification sequestrate carbon dioxide, co2, greenhouse gases
peak shaving energy management, energy storage sequestration carbon dioxide, co2, greenhouse gases
pemfc proton exchange membrane fuel cell, hydrogen fuel cell simultaneously practice thrift resource efficiency, waste reduction
pervaporation membrane separation, desalination smart grid renewable energy integration, energy efficiency
photocatalyst - smart grids renewable energy integration, energy efficiency
photocatalytic - sncr denitration flue gas treatment, air pollution control
photochemistry solar energy conversion sofc solid oxide fuel cell, clean energy
photodegradable - soilless cultivation -
photolysis - soilless culture -
photolytic - solar energy, cell, panel, power
photooxidation - solarenergy -
photoreaction solar energy conversion solarpanel -
photosynthetic biofuels solarpower -
photovoltaic - straw biomass resource, renewable energy
pollutant reduce, reduction, capture, remove, removal sugar beet biofuel feedstock, renewable energy
polluted clean, filter sun shading sustainable design, building efficiency
pollution reduce, reduction, capture, remove, removal, avoid sunshade sustainable design, building efficiency
polution reduce, reduction, capture, remove, removal, avoid supercapacitor energy storage, renewable energy integration
polylactide - supercharge electric vehicle, battery
poultry farming sustainable agriculture, waste management supercharging electric vehicle, battery
poultry manure sustainable agriculture, waste management sustainability goal, practice, system, production
propeller electric, hybrid, low-emission sustainable system, production, process, technology, resource, energy
propulsion electric, hybrid, low-emission thermocatalytic biomass conversion, waste valorization
propulsor electric, hybrid, low-emission thermochemical biomass conversion, renewable energy
pyrohydrolysis waste conversion, biomass conversion ultrafilter water treatment, purification
recuperation energy, waste, material, resource ultrafiltrate water treatment, purification
recyclable energy, waste ultrafiltration water treatment, purification
recyclable material - ultralight material, construction
recyclate energy, waste, material, resource ultralow emission, power consumption
recycle energy, waste, material, resource urban planning sustainable development, green infrastructure
recycle material - uv photolysis water treatment, pollutant degradation
recycle practice thrift - waste management recycling, resource recovery
recycle product - waste recycle -
recycleable - wasted recycling, resource recovery, avoid
recycled energy, waste, resource wastepaper recycling, paper recycling, avoid
recycling energy, waste, material, resource water purification -
regasification liquefied natural gas, renewable energy sources water treatment low energy, reuse, recycle, efficient
regenerate soil, environment, material, resource waterpower -
regenerated soil, environment, material, resource windmill -
regeneration soil, environment, material, resource windpower -
regenerative braking, energy recovery windturbine -
remediation environmental cleanup, contaminated soil windwheel -
renewability energy, waste, material, resource woody biomass, biofuel
renewable energy, resource, material
  • Note: The table reports the list of keywords for REGEX match into patents texts produced with the W2V model, i.e. a text containing at least one of the expression combinations, if classified as “green” by standard classification algorithms, were classified as “true green”.
@@ A.3. Word2Vec specification
In this section, we provide details about the selection of the Word2Vec model’s hyperparameters. We need to find the context window size ($C$), which defines the context length for training the neural network, and the minimum count ($MC$), which sets the frequency threshold for a word to be included in the training corpus. We employ a data-driven tuning process. We use WordSim353, which is a test collection for measuring word similarity or relatedness (Agirre et al., 2009). We then train multiple Word2Vec models using a grid of different $W$ and $MC$ parameter combinations. For each trained model, we calculated the cosine similarity for word pairs in the WordSim353 list and measured the Pearson correlation between our results and the gold standard scores. As shown in Figure A1, all tested parameter pairs yielded a positive correlation above 50%. Generally, models with a smaller $C$ performed better when the minimum count was held constant. This empirical process pointed towards optimal values of approximately MCMC=40 and CC=2, which we adopt in our analysis in the manuscript.
@@ B.1. Firm-level dataset distribution
01-03 Crop, animal production, fishing, hunting, forestry and logging 79,028 2.95
05-09 Mining and extraction 5,914 0.22
10-33 Manifacture 335,353 12.53
35 Electricity, gas, steam and air conditioning supply 24,245 0.91
36-39 Water, sewerage and waste management 15,266 0.57
41-43 Construction, civil engineering and special activities 297,370 11.11
45-47 Wholesale and retail trade 630,142 23.54
49-53 Transport and storage 126,121 4.71
55-56 Accommodation and food service activities 133,306 4.98
58-63 Information and communication 110,921 4.14
64-69 Financial, insurance and real estate activities 284,155 10.62
69-75 Professional, scientific and technical activities 311,900 11.65
77-79 Administrative and support service activities 39,906 1.49
80-82 Security and investigation activities 68,289 2.55
@@
85 Education 32,992 1.23
86-88 Human health and social work activities 89,969 3.36
90-93 Arts, entertainment and recreation 39,329 1.47
94-96 Other services activities 51,716 1.93
97-98 Activities of households as employers; undifferentiated goods 30 <1<1
99 Activities of extraterritorial organisations and bodies 51 <1<1
@@
Austria 385 147 26
Belgium 529 179 33
Bulgaria 206 18 1
@@
Spain 1,200 232 39
Finland 813 213 33
France 2,681 668 148
Greece 16 3 0
Croatia 20 3 0
Hungary 108 25 3
@@
Netherlands 187 63 13
Poland 984 228 59
Portugal 125 30 3
Romania 146 44 21
Sweden 1,156 271 36
Slovenia 86 19 4
@@
01-03 Crop, animal production, fishing, hunting, forestry and logging 56 12 1